OpenStudy (anonymous):

Which of the following is the slope between the points (-6, 0) and (2, -9)? eight ninths -negative eight ninths nine eighths -negative nine eighths

OpenStudy (karatechopper):

y1-y2/x1-x2

OpenStudy (karatechopper):

0+9/-6-2 9/-8 -9/8

OpenStudy (karatechopper):

negative nine eighths

OpenStudy (karatechopper):

please click good answer

OpenStudy (anonymous):

slope m=(y2-y1)/(x2-x1)=-9/8